In a chemistry lab the concentration of one solution is 10% HCU and that of a second solution is 60% HCI. How many milliliters of each should be mixed to obtain 50 ml of a 30% solution?
Answer by robertb(5830)

0.10x + 0.60(50 - x) = 0.30*50
0.10x + 30 - 0.60x = 15,
-0.5x = -15,
x = 30, #ml of 10% solution,
50 - 30 = 20, #ml of 60% solution.